OpenStudy (anonymous):

The point (3, 2) is 6sqrt(2) units away from the point (x, 8). Find the value of x

OpenStudy (s3a):

d^2 = (x2 - x1)^2 + (y2-y1)^2 6^2 * sqrt(2)^2 = (x - 3)^2 + (8-2)^2 36*2 = (x-3)^2 + 6^2 36*2 = (x-3)^2 + 36 36*2 - 36 = (x-3)^2 + 36 - 36 36 = (x-3)^2 6 = x-3 6 + 3 = x - 3 + 3 9 = x x = 9

OpenStudy (anonymous):

I have the answers and it says x=-3

OpenStudy (anonymous):

I just dont know how to get it

OpenStudy (s3a):

Actually, my x = 9 works but so does x = -3 and I did skip something accidentally but I found it and it's that when I square rooted both sides 36 = (x-3)^2 I should have taken both the positive and negative possibilities. sqrt(36) = +/- sqrt((x-3)^2) so 6 = +(x-3) and 6 = -(x-3) so x = 9 AND x = -3.

OpenStudy (anonymous):

Oh i see now. Thats actually pretty easy, just didnt know where to start.

OpenStudy (s3a):

I should say "or" rather than "and".

OpenStudy (s3a):

Yes, well I'm glad you get it now. :)

OpenStudy (anonymous):

Thank you =D

OpenStudy (s3a):

No problem!